### v1

# Galois theory, automorphism groups of prime models, and the Picard-Vessiot closure

Abstract
The main result of the paper extends the well-known Galois correspondence between closed subgroups of $Aut(acl^{eq}(A)/A)$ and dcldcl-closed intermediate sets $A\subseteq C\subseteq acl^{eq}(A)$ to the setting of the minimal closure inside of a prime model of a totally transcendental theory $T=T^{eq}$, $A\subseteq mcl^{eq}(A)\subseteq M_{A}$ (namely, $mcl^{eq}(A)$ is the intersection of all self elementary embeddings of the prime model $M_{A}$). Precisely, for $B$ a normal intermediate extension $A\subseteq B\subseteq mcl^{eq}(A)$, definable subgroups of $Aut(B/A)$ are in Galois correspondence with dcldcl-closed intermediate sets $A\subseteq C\subseteq B$.
This applies in particular to the “Picard-Vessiot closure” $K^{PV_{\infty}}$ (or $K_{\infty}$) of a differential field $K$ of characteristic 0 with algebraically closed field $C_{K}$ of constants. We also show that normal differential subfields of $K^{PV_{\infty}}$ containing $K$ are “iterated $PV$-extensions” of $K$, and the Galois correspondence result above holds for these extensions. This fills in some missing parts of Magid’s paper [5].
We also discuss exact sequences $1\to N\to G\to H\to 1$, where $G=Aut(K_{2}/K)$, $N=Aut(K_{2}/K_{1})$ and $H=Aut(K_{1}/K)$, $K_{1}$ is a (maybe infinite type) $PV$ extension of $K$. $K_{2}$ is a (maybe infinite type) $PV$ extension of $K_{1}$ and $K_{2}$ is normal over $K$ (in the differential closure of $K$) and again $C_{K}$ is algebraically closed. Both $N$ and $H$ have the structure of proalgebraic groups over $C_{K}$. We show that conjugation by any given element of $G$ is a proalgebraic automorphism of $N$. Moreover if $G$ splits as a semidirect product $N\rtimes H$, then left multiplication by any fixed element of $G$ is a morphism of proalgebraic varieties $N\times H\to N\times H$. This improves and extends observations in Section 4 of [5] which dealt with one particular example.
1 Introduction
In this paper we work in the context of a totally transcendental theory $T$ eliminating imaginaries ($T=T^{eq}$). $\bar{M}$ will denote a very saturated model, which exists as $T$ is totally transcendental ($t.t.$). Typically, $A$ and $B$ will denote small definably closed subsets of $\bar{M}$. For $A\subseteq B$, $Aut(B/A)$ denotes the group of $A$-elementary permutations of $B$. In the context of a $t.t.$ theory, one has, over arbitrary parameter sets $A$, prime models $M_{A}$ which are unique up to $A$-isomorphism and which are atomic and strongly $\omega$-homogeneous over $A$ (whenever $\bar{a}$, and $\bar{b}$ are finite tuples in $M_{A}$ with the same type over $A$, there is an automorphism $\sigma\in Aut(M_{A}/A)$ taking $\bar{a}$ to $\bar{b}$).
There are two main notions used in this paper to set up the Galois correspondence. Firstly, is that of a (relatively) definable subset of an automorphism group of a prime model. These were defined in [9] and used to show existence and uniqueness of invariant measures on automorphism groups under certain assumptions. In [9], this notion was used to give a characterization of definable cocycles (first defined in [8]). This was extended in [7]. Also in [9] this notion is related to differential Galois theory.
The second notion used to set up the correspondence is the minimal closure of a set of parameters $A$ contained in the prime model $M_{A}$, which we denote $mcl(A)$. The minimal closure $mcl(A)$ is the intersection of the images all elementary embeddings of the prime model $M_{A}$ into itself.
In [13], $mcl(A)$ is defined and strong homogeneity and closure properties of $mcl(A)$ are shown in anticipation of [12] where they were used to give the first model-theoretic description of the Galois correspondence for the Picard-Vessiot, and strongly normal differential Galois theories. The other major result of [12] is the well-known “infinite” Galois correspondence for aclacl, developed using the profinite topology on the automorphism group. However, in [12], working in $DCF_{0}$, the theory of differential fields of characteristic 0, (essentially the context of a prime model of a $t.t.$ theory) only the “finite type” Picard-Vessiot extensions are studied. No topology or definable structure is given to the automorphism groups of infinitely dcldcl-generated extensions. The structure on the automorphism groups associated to finitely dcldcl-generated extensions is given by the binding group construction.
In this paper, we replace the role of the topology of pointwise convergence (or the profinite topology) on $Aut(acl^{eq}(A)/A)$ with the notion from [9] of a definable group of automorphisms to give a satisfying “infinite” Galois theory for $Aut(mcl^{eq}(A)/A)$.
Before stating this main theorem we give a few more definitions and say something about our approach. Let $A$ be a parameter set and $M_{A}$ be a copy of the prime model over $A$. Then an intermediate set $A\subseteq B\subseteq M_{A}$ is said to be normal (in $M_{A}$ over $A$) if $B$ is fixed setwise by $Aut(M_{A}/A)$ in which case $Aut(M_{A}/A)$ has $Aut(M_{A}/B)$ as a normal subgroup with a surjective quotient map to $Aut(B/A)$ given by restriction. Next, we call $B$ minimal (in $M_{A}$ over $A$) if there is no $C$ with $A\subseteq C\subseteq B$ such that $C$ is Tarski-Vaught in $B$. We show that, $mcl(A)$, is the largest (under inclusion) normal minimal subset of $M_{A}$ over $A$, i.e. any normal minimal subset of $M_{A}$ over $A$ is contained in the minimal closure and of course any normal subset $B$ of $mcl(A)$ containing $A$ must be minimal.
The main result is:
Theorem 1.1.
For $B$, a normal and minimal set in $M_{A}$ over $A$, equivalently an intermediate extension $A\subseteq B\subseteq mcl^{eq}(A)$, definable subgroups of $Aut(B/A)$ correspond to dcldcl-closed intermediate sets $A\subseteq C\subseteq B$.
In fact, our statement in Section 2 (Theorem 2.13) will state this correspondence for normal minimal sets without explicit reference to the minimal closure; the reason being that we find it more comfortable to work with the notion of Tarski-Vaught minimality than with the minimal closure, although we end section 2 with the necessary comparison of notions.
We now discuss the example of the theory $DCF_{0}$ of differentially closed fields of characteristic zero and Picard-Vessiot extensions of a differential subfield. $DCF_{0}$ is $\omega$-stable so the paragraphs above apply. The prime model over a differential subfield $K$ is called the differential closure of $K$ and often denoted $K^{diff}$. The classical notion of a Picard-Vessiot extension of $K$ (assuming the field $C_{K}$ of constants is algebraically closed) is a differential field extension of $K$ generated over $K$ by a “fundamental system of solutions” in $K^{diff}$ of a linear differential equation $\partial Y=AY$ over $K$, namely $Y$ is a column vector of unknowns and $A$ an $n\times n$ matrix over $K$. It is convenient to follow the conventions of Wibmer in for example [14], where this is referred to as a Picard-Vessiot extension for the equation $\partial Y=AY$ . More generally Wibmer defines a Picard-Vessiot extension of $K$ for a (possibly infinite) family $\mathcal{F}$} of linear differential equations over $K$ as the differential field generated over $K$ by the union of all fundamental systems of solutions (in $K^{diff}$) of all equations in $\mathcal{F}$}. He calls a Picard-Vessiot extension “of finite type” if it is the Picard-Vessiot extension for a single linear differential equation (equivalently finitely many). With these conventions there is a greatest Picard-Vessiot extension of $K$ (inside a given copy of $K^{diff}$), which we call $K^{PV}$. Then $K^{PV}$ will be normal and minimal over $K$ and $Aut(K^{PV}/K)$ has the structure of a proalgebraic group (with respect to $C_{K}$). The closed subgroups in our case correspond to Zariski closed subgroups and Theorem 1.1 gives the well-known Galois correspondence. Actually $K^{PV}$ also has the structure of a prodefinable or pro-differential algebraic group defined over $K$, which is isomorphic over $K^{PV}$ to the corresponding proalgebraic over $C_{K}$ group.
In any case Theorem 1.1 covers other cases where it gives new results. One such is the “Picard-Vessiot closure” of a differential field which will be looked at in Section 3. Given again a differential field $K$ with algebraically closed field of constants, we have as above $K^{PV}$. In contradistiction with the case of $K^{alg}$ the algebraic closure of $K$, we typically have that $(K^{PV})^{PV}$ properly contains $K^{PV}$. For example when $K=\mathbb{C}$, the complex field considered as a field of constants, then $Aut(K^{PV}/K)$ is the connected component of the proalgebraic group on one generator. On the other hand $\mathbb{C}(t)$ (with $d/dt$) is contained in $K^{PV}$ and $Aut(\mathbb{C}(t)^{PV}/\mathbb{C}(t))$ is the free proalgebraic group on continuum many generators (and of course is contained in $Aut((K^{PV})^{PV}/K^{PV})$). We define $K^{PV_{0}}=K$, $K^{PV_{n+1}}=(K^{PV_{n}})^{PV}$, and $K^{PV_{\infty}}=\cup_{n}K^{PV_{n}}$.
Already in [12] and [13] where the minimal closure is introduced, it is stated that $K^{PV}$ is normal and minimal being contained in the minimal closure $K\subseteq K^{PV}\subseteq mcl(K)\subseteq K^{diff}$. But in fact, more is true:
Proposition 1.2.
Each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ are normal and minimal over $K$ in $K^{diff}$, i.e. $K\subseteq K^{PV}\subseteq(K^{PV})^{PV}\subseteq...\subseteq K^{PV_{\infty}}\subseteq mcl(K)\subseteq K^{diff}$.
Hence Theorem 1.1 applies to give the Galois correspondence. Magid calls $K^{PV_{\infty}}$ the complete PV closure of $K$ (although $PV$ closure would be correct and enough). In [5] a full Galois correspondence for $Aut(K^{PV_{\infty}}/K)$ is given; between intermediate differential fields and “closed” subgroups. However no definition of closed subgroup is offered (other than the tautological definition of being precisely the subgroups of the form $Fix(L)$ for $L$ an intermediate differential field). So Theorem 1.1 fills this gap in a reasonably canonical form. Neverthess there is some content in [5], namely that the set of fixed points of $Fix(L)$ is precisely $L$. The same observation, in a more general model-theoretic environment, is made in [12] ( top of p. 1166).
Actually Magid’s Galois correspondence (Theorem 5 of [5]) is for what he calls normal “locally iterated” $PV$ extensions. By our conventions that $PV$ extensions may be of infinite type, we could and will rename these “iterated $PV$ extensions”. Such an object is a differential field $L$ with $K\leq L\leq K^{PV_{\infty}}$ such that EITHER $L$ is the union of a strictly increasing chain $(L_{i}:i<\omega)$, with $L_{0}=K$ and each $L_{i+1}$ a $PV$ extension of $L_{i}$, OR there is a finite such chain $(L_{i}:i<n)$ with $L=L_{n}$. In particular each $K^{PV_{n}}$ as well as $K^{PV_{\infty}}$ are examples of iterated $PV$ extensions.
We will prove:
Proposition 1.3.
Let $K\leq L\leq K^{PV_{\infty}}$ be normal over $K$. Then $L$ is an iterated $PV$ extension of $K$ where moreover each $L_{i}$ is also normal and minimal (over $K$ in $K^{diff}$).
Hence Theorem 1.1 applies to any normal $L\leq K^{PV_{\infty}}$ containing $K$.
Now for $n>1$, $Aut(K^{PV_{n}}/K)$ can not have any structure as a prodefinable group in $DCF_{0}$ or equivalently as a proalgebraic group with respect to the constants. So one can ask what additional information (other than having a composition series of proalgebraic groups in the constants) does one have.
We will do an analysis of the case where $n=2$. In fact we work at the more general level of $2$-iterated PVPV-extensions. Let $K_{1}$ be a PV extension of $K$, and let $K_{2}$ be a PV extension of $K_{1}$ such that $K_{2}$ is normal in $K^{diff}$ over $K$. Both $K_{1}$ and $K_{2}$ may be of infinite type. A special case would be $K_{1}=K^{PV_{1}}$ and $K_{2}=K^{PV_{2}}$.
We let $G=Aut(K_{2}/K)$, $N=Aut(K_{2}/K_{1})$, and $H=Aut(K_{1}/K)$. Then $N$ and $H$ have the structure of proalgebraic groups with respect to the algebraically closed field $C_{K}$, and we have the exact sequence
$1\to N\xrightarrow{i}G\xrightarrow{\pi}H\to 1$
where $i$ is inclusion.
Proposition 1.4.
(i) Let $\tau\in G$. Then conjugation by $\tau$ is a proalgebraic automorphism of $N$.
(ii) Suppose that $G$ splits as the semidirect product of $N$ and $H$ then left translation by a given element of $G$ is a proalgebraic morphism from the underlying proalgebraic variety of $G$, $N\times H$, to itself.
The proof of Proposition 1.4 will go through the the general structure theory developed in [1].
In Section 4 of [5] Magid studies a particular example. The base field $K$ is ${\mathbb{C}}(t)$. $F_{1}$ is a certain finite type Picard-Vessiot extension of $K$ with Galois group $H$ equal to the additive group of the constants. $E$ is a certain (infinite type) Picard-Vessiot extension of $F_{1}$ which is normal (in $K^{diff}$ over $K$) and whose Galois group $N$ is the direct product of continuum many copies of the additive group of the constants, indexed by the constants, a proalgebraic group. Magid observes that if $G=Aut(E/K)$ then $G$ is the semidirect product of $N$ and $H$. In this special case Magid computes that left multiplication by any given element of $G$ is a proalgebraic morphism of the proalgebraic variety $N\times H$ to itself (but not right multiplication). Of course this is just a special case of our Proposition 1.4 (ii).
In passing, we know of no (nontrivial) example where $Aut(K^{PV_{2}}/K)$ is a semidirect product of $Aut(K^{PV_{2}}/K^{PV_{1}})$ and $Aut(K^{PV_{1}}/K)$.
Acknowledgements. The second author would like to thank Andy Magid and Bruno Poizat for discussions and questions, especially concerning the automorphism group of $K^{PV_{2}}$ over $K$.
2 The Galois theory
We will here prove Theorem 1, after expanding on the notions introduced in the introduction above.
We take $T$ to be a complete $t.t.$ theory in language $L$. When $L$ is countable this means precisely that $T$ is $\omega$-stable. As in Section 1, we work in a saturated monster model ${\bar{M}}^{eq}$, $A$ and $B$ denote small definably closed subsets, and $M_{A}$ denotes (a copy of) the prime model over $A$. We fix $B$ with $A\subseteq B\subseteq M_{A}$. We defined $B$ to be normal (in $M_{A}$ over $A$) if $B$ is setwise invariant under $Aut(M_{A}/A)$. This just means that for any finite tuple $b$ from $B$, all realizations of $tp(b/A)$ in $M_{A}$ are already in $B$. By $Aut(B/A)$ we denote the group of elementary (in the sense of $M_{A}$ or $\bar{M}$}) over $A$ permutations of $B$. By our assumptions this is precisely the set of restrictions to $B$ of elements of $Aut(M_{A}/A)$. In [9] we introduced “definable” subsets of $Aut(B/A)$, which we will now call relatively definable subsets (reverting to the original language from [2]): For a formula $\phi({\bar{x}},{\bar{y}})$ over $A$, with $\bar{x}$}, $\bar{y}$} finite tuples of variables of the same length, and a tuple $\bar{b}$} from $B$ of the same length, $\{\sigma\in Aut(B/A):\models\phi(\sigma({\bar{b}}),{\bar{b}})\}$ is a relatively definable subset of $Aut(B/A)$.
We now let $\bar{b}$} be an enumeration of $B$, and $p_{0}({\bar{x}})=tp({\bar{b}}/A)$. So now $\bar{x}$} is a possibly infinite tuple of variables. We let $S_{p_{0}}(B)$ be the (Stone) space of complete types $p({\bar{x}})$ over $B$ which extend $p_{0}$. In [9] we pointed out that relatively definable subsets of $Aut(B/A)$ are in natural one-one-correspondence with clopen subsets of $S_{p_{0}}(B)$. (We will repeat this below). It is natural to ask: What corresponds to arbitrary closed subsets of $S_{p_{0}}(B)$?
Definition 2.1.
Let $X$ be a closed subset of $S_{p_{0}}(B)$ given by the set $\Sigma({\bar{x}},{\bar{b}})$ of formulas, where $\Sigma({\bar{x}},{\bar{y}})$ is a collection of $L_{A}$-formulas and $\bar{b}$} an enumeration of $B$.
(i) $X(Aut({\bar{M}}/A))=\{\sigma\in Aut({\bar{M}}/A):{\bar{M}}\models\Sigma(\sigma({\bar{b}}),{\bar{b}})\}$, and
(ii) $X(Aut(B/A))=X(Aut({\bar{M}}/A))|_{B}$ $(=\{\sigma\in Aut(B/A):\models\Sigma(\sigma({\bar{b}}),{\bar{b}})\})$.
Lemma 2.2.
(i) Let $X$ be a closed subset of $S_{p_{0}}(B)$. Then $X$ is determined by $X(Aut({\bar{M}}/A))$.
(ii) Let $X$ be a clopen subset of $S_{p_{0}}(B)$. Then $X$ is determined by $X(Aut(B/A))$.
Proof.
(i) Suppose $X,Y$ are closed subsets of $S_{p_{0}}(B)$ given by $\Sigma({\bar{x}},{\bar{b}})$ and $\Gamma({\bar{x}},{\bar{b}})$ respectively. Suppose $X(Aut({\bar{M}}/A))=Y(Aut({\bar{M}}/A))$. Let $\bar{c}$} realize $\Sigma({\bar{x}},{\bar{b}})$. As $\bar{c}$} realizes $p_{0}$, there is $\sigma\in Aut({\bar{M}}/A)$ with $\sigma({\bar{b}})={\bar{c}}$. So $\sigma\in X(Aut({\bar{M}}/A))$, whereby $\sigma\in Y(Aut({\bar{M}}/A))$, whereby $\bar{c}$} realizes $\Gamma({\bar{x}},{\bar{b}})$. This suffices.
(ii) This is essentially Proposition 2.8 (ii) of [9]. But as our notation here is slightly different, we give the proof. We let $x,y$ denote finite tuples of variables of an appropriate length. Let $b_{0}$ be a finite tuple from $B$ and let $\phi(x,y)$, $\psi(x,y)$ be formulas over $A$, such that for $\sigma\in Aut(B/A)$, $\models\phi(\sigma(b_{0}),b_{0})$ iff $\models\psi(\sigma(b_{0}),b_{0})$. We want to prove that $\phi(x,b_{0})$ is equivalent to $\psi(x,b_{0})$ in $S_{p_{0}}(B)$. Let $\chi(y)$ isolate $tp(b_{0}/A)$. We may assume that each of $\phi(x,b_{0})$, $\psi(x,b_{0})$ implies $\chi(x)$. We will show that $\phi(x,b_{0})$ and $\psi(x,b_{0})$ are equivalent. For this it suffices to work in the model $M_{A}$. If the formulas are not equivalent then without loss there is $c_{0}\in M_{A}$ such that $\models\phi(c_{0},b_{0})\wedge\neg\psi(c_{0},b_{0})$. As $tp(c_{0}/A)=tp(b_{0}/A)$, there is $\sigma\in Aut(M_{A}/A)$ such that $\sigma(b_{0})=c_{0}$. Now $c_{0}\in B$ by normality and $\tau=\sigma|B\in Aut(B/A)$. So $\models\phi(\tau(b_{0}),b_{0})$ but $\models\neg\psi(\tau(b_{0}),b_{0})$, a contradiction. ∎
Remark 2.3.
Clearly in general an arbitrary closed subset $X$ of $S_{p_{0}}(B)$ need not be determined by $X(Aut(B/A))$. For example suppose $b_{i}\in B$ is not in $acl(A)$ and let $\Sigma({\bar{x}},{\bar{b}})$ express that $x_{i}\neq b_{j}$ for all $j$. Then $\Sigma({\bar{x}},{\bar{b}})$ determines a nonempty closed subset of $S_{p_{0}}(B)$, but there is no $\sigma\in Aut(B/A)$ such that $\models\Sigma(\sigma({\bar{b}}),{\bar{b}}))$.
We get a Stone topology on $Aut({\bar{M}}/A)$ induced from $S_{p_{0}}(B)$ and we call the corresponding subsets of $Aut({\bar{M}}/A)$, $B$-closed. It then makes sense to talk about a $B$-closed subgroup of $Aut({\bar{M}}/A)$.
Definition 2.4.
By a closed subgroup $G$ of $Aut(B/A)$ we mean $H|_{B}$ for some $B$-closed subgroup $H$ of $Aut({\bar{M}}/A)$. We will say that $\Sigma({\bar{x}},{\bar{b}})$ defines $G$ if it defines $H$ in $Aut({\bar{M}}/A)$.
Note that a closed subgroup of $Aut(B/A)$ is in particular a subgroup of $Aut(B/A)$ which is of the form $X(Aut(B/A))$ for some closed subset of $S_{p_{0}}(B)$, but this is NOT an equivalent description.
The next Lemma holds with no conditions on $T$, only that $A\subseteq B$ are small subsets of $\bar{M}$} (and Definition 2.1 (i) makes sense at this general level).
Lemma 2.5.
$\Sigma({\bar{x}},{\bar{b}})$ defines a $B$-closed subgroup $H$ of $Aut({\bar{M}}/A)$ if and only if $\Sigma({\bar{x}},{\bar{y}})$ defines an equivalence relation on realizations of $p_{0}$ (in $\bar{M}$}).
Proof.
This is the same as the proof of Lemma 3.1 of [9] which is about the case when $X$ is clopen (so given by a single formula) and $B=M_{A}$. We will give it anyway. First prove left implies right. So we assume $\Sigma({\bar{x}},{\bar{b}})$ defines a $B$-closed subgroup $H$ of $Aut({\bar{M}}/A)$
(i) Reflexive: As the identity is in $H$, $\Sigma({\bar{b}},{\bar{b}})$ holds whereby $\Sigma({\bar{x}},{\bar{x}})\subseteq p_{0}({\bar{x}})$.
(ii) Symmetric: Suppose that $\bar{c}$} realizes $p_{0}$ in $\bar{M}$ and $\models\Sigma({\bar{c}},{\bar{b}})$. Then there is $\sigma\in H$ such that $\sigma({\bar{b}})={\bar{c}}$. But then $\sigma^{-1}\in H$, whereby $\models\Sigma(\sigma^{-1}({\bar{b}}),{\bar{b}})$. As $\Sigma({\bar{x}},{\bar{y}})$ is over $A$ we can apply $\sigma$ to get $\models\Sigma({\bar{b}},\sigma({\bar{b}})$, namely $\models\Sigma({\bar{b}},{\bar{c}})$.
We have shown that $p_{0}({\bar{x}})$ and $\Sigma({\bar{x}},{\bar{b}})$ implies $\Sigma({\bar{b}},{\bar{x}})$. Hence $\Sigma$ is symmetric on all realizations of $p_{0}$.
(iii) Transitive: Given that we have symmetry, it suffices to prove that if $\bar{c}$}, $\bar{d}$} realize $p_{0}$ and $\models\Sigma({\bar{c}},{\bar{b}})$ and $\models\Sigma({\bar{d}},{\bar{b}})$, then $\models\Sigma({\bar{d}},{\bar{c}})$. So we have $\sigma,\tau\in H$ such that $\sigma({\bar{b}})={\bar{c}})$ and $\tau({\bar{b}})={\bar{d}}$. Then $\sigma^{-1}\tau\in H$, whereby $\models\Sigma(\sigma^{-1}\tau({\bar{b}}),{\bar{b}})$, so $\models\Sigma(\sigma^{-1}({\bar{d}}),{\bar{b}})$. Apply $\sigma$ to get $\Sigma({\bar{d}},{\bar{c}})$, as required.
The right implies left direction just reverses the proofs above. For example if $\Sigma({\bar{x}},{\bar{x}})$ holds for $\bar{x}$} realizing $p_{0}$ then $\Sigma({\bar{b}},{\bar{b}})$ holds, so the identity is in $H$. ∎
We now bring in (some of) the assumptions.
Lemma 2.6.
$\Sigma({\bar{x}},{\bar{y}})$ be as in Lemma 2.5. Then there are $A$-definable equivalence relations $\phi_{i}({\bar{x}},{\bar{y}})$ such that on realizations of $p_{0}$, $\Sigma({\bar{x}},{\bar{y}})$ agrees with $\{\phi_{i}({\bar{x}},{\bar{y}}):i\in I\}$.
Proof.
This is precisely elimination of hyperimaginaries in stable theories. See for example [10] or [3]. ∎
Proposition 2.7.
(i) The $B$-closed subgroups of $Aut({\bar{M}}/A)$ are precisely subgroups of the form $\{\sigma\in Aut({\bar{M}}/A):\sigma$ fixes $C$ pointwise} for some $C\subseteq B$.
(i) Likewise the closed subgroups of $Aut(B/A)$ are precisely subgroups of the form $\{\sigma\in Aut(B/A):\sigma$ fixes $C$ pointwise
} for some $C\subseteq B$.
Proof.
Remember we are working in ${\bar{M}}^{eq}$. First fix a subset $C$ of $B$. Suppose $c\in C$ and suppose $c=b_{\alpha}$ with respect to the given enumeration $\bar{b}$} of $B$. Then the formula $x_{\alpha}=b_{\alpha}$ defines $\{\sigma\in Aut({\bar{M}}/A):\sigma(c)=c\}$. So the collection of all these formulas, as $c$ ranges over $C$, defines the $B$-closed subgroup of $Aut({\bar{M}}/A)$ which is precisely the set of $\sigma$ which fix $C$ pointwise.
Conversely, let $H$ be a closed subgroup of $Aut({\bar{M}}/A)$. By Lemma 2.7 it is defined by a collection of formulas $\phi_{i}({\bar{x}},{\bar{b}})$ ($i\in I$) where each $\phi({\bar{x}},{\bar{y}})$ is a formula over $A$ defining an equivalence relation on realizations of $p_{0}$. Fix $i$. Then only finitely many variables, actually appear in $\phi_{i}$ which we assume for simplicity to be of the form $(x_{1},..,x_{n},y_{1},..,y_{n})$ (where $y_{1},..,y_{n}$ correspond to $b_{1},..,b_{n}$). Then $\phi_{i}$ defines an equivalence relation $E_{i}$ on realizations of $tp(b_{1},..,b_{n}/A)$, and $\phi_{1}(x_{1},..,x_{n},b_{1},..,b_{n})$ defines $Fix(b_{1},..,b_{n}/E_{i})$, the automorphisms which fix this imaginary $c_{i}=(b_{1},..,b_{n})/E_{i}$. But $B$ is definably closed in ${\bar{M}}^{eq}$, so $c_{i}\in B$. Let $C\subseteq B$ be the set of these $c_{i}$.
(ii) follows as the closed subgroups of $Aut(B/A)$ are by definition the intersections of $B$-closed subgroups of $Aut({\bar{M}}/A)$ with $Aut(B/A)$. ∎
Remark 2.8.
(i) Proposition 2.7 is valid for any definably closed $B$ (containing $A$) under the additional elimination of hyperimaginaries assumption. Namely there is no prime model $M_{A}$ in the picture.
(ii) From Proposition 2.7 we get a tautological Galois correspondence between $B$-closed subgroups of $Aut({\bar{M}}/A)$ and definably closed subsets of $B$ containing $A$. All that has to be shown is that if $C$ is a subset of $B$ containing $A$, then the set of elements of $B$ (even of $\bar{M}$}) fixed by the group of automorphisms of $\bar{M}$ which fix $C$ pointwise is $dcl(C)$, which is immediate by saturation and homogeneity of $\bar{M}$.
(iii) In the context of $T$ being $t.t.$, and $B$ normal over $A$ in $M_{A}$, obtaining a full Galois correspondence between closed subgroups of $Aut(B/A)$ and definably closed subsets of $B$ containing $A$, depends again on showing that the elements of $B$ fixed by the set (group) of $\sigma\in Aut(B/A)$ which fix some $C\subseteq B$ is precisely $dcl(A,C)$. This requires an additional assumption on $B$ which we introduce below.
Recall the basic notion of “Tarski-Vaught” (T-V) subset: Let $C\subseteq D$ be sets of parameters in $\bar{M}$. We will say that $C\leq_{T-V}D$ (in the sense of $\bar{M}$) if whenever $\phi(x)$ is an $L_{C}$-formula such that ${\bar{M}}\models\phi(d)$ for some $d$ from $D$ then there is $c$ from $C$ such that ${\bar{M}}\models\phi(c)$.
Our underlying assumptions are still that $T=T^{eq}$ is $t.t.$ $A\subseteq B$ are definably closed sets, $B\subseteq M_{A}$ and $B$ is normal over $A$ in $M_{A}$.
Definition 2.9.
We say that $B$ is minimal (over $A$ in $M_{A}$) if there is no $C\supseteq A$ with $C\leq_{TV}B$ and $C\neq B$.
Remark 2.10.
We know (see Lemma 18.4 of [11] for example) that normality of $B$ over $A$ (in $M_{A}$) implies that $M_{A}$ is also prime over $B$. The following gives an extension.
Lemma 2.11.
Suppose $B$ is (normal and) minimal over $A$. Then for any $C$ with $A\subseteq C\subseteq B$, $M_{A}$ is also prime over $C$.
Proof.
Let $M_{C}$ be a prime model over $C$ contained in $M_{A}$.
Claim. $B\subseteq M_{C}$.
Proof of claim. Consider $D=M_{C}\cap B$. Let $d$ be a (finite) tuple from $D$ and $b^{\prime}$ a (finite) tuple from $B$ and $\phi$ a formula over $A$ say such that $\models\phi(d,b^{\prime})$. Now $tp(b^{\prime}/A)$ is isolated by a formula $\psi(y)$. So the formula $\phi(d,y)\wedge\psi(y)$ is realized in $M_{A}$, so also realized in $M_{C}\prec M_{A}$, by $b^{\prime\prime}$ say. But all realizations of $\psi(y)$ in $M_{A}$ are in $B$ (by normality), so $b^{\prime\prime}\in M_{C}\cap B$. We have shown that $D\leq_{TV}B$. Hence by our assumptions $D=B$ and so $B\subseteq M_{C}$. ∎
Now $M_{C}$ is clearly atomic over $A$, so prime over $A$. Now (by the claim) $B\subseteq M_{C}$ and is clearly normal over $C$ in $M_{C}$ whereby, using Remark 2.11, $M_{C}$ is also prime over $B$. So by Remark 2.11 and uniqueness $M_{C}$ is isomorphic to $M_{A}$ over $B$, so also over $C$, and so $M_{A}$ is prime over $C$. ∎
Lemma 2.12.
Assume $B$ (normal and) minimal in $M_{A}$. Let $C$ be definably closed with $A\subseteq C\subseteq B$. Let $H=\{\sigma\in Aut(B/A):\sigma$ fixes $C$ pointwise}. Then $C$ is precisely the set of elements of $B$ fixed by $H$.
Proof.
Suppose that $b\in B\setminus C$. As $M_{A}$ is prime over $C$, $tp(b/C)$ is isolated. As $b\notin dcl(C)$, $tp(b/C)$ has another realization $b^{\prime}$ in $M_{A}$ which will be in $B$ by normality., As $M_{A}$ is $\omega$-homogeneous over $C$ there is $\sigma\in Aut(M_{A}/C)$ with $\sigma(b)=b^{\prime}$. Then $\sigma|B\in Aut(B/A)$ and moves $b$.
Hence, bearing in mind Proposition 2.7 (ii), we have:
Theorem 2.13.
Suppose $A\subseteq B\subseteq M_{A}$ with $B$ definably closed and both normal and minimal (over $A$ in $M_{A}$). Then there is a Galois correspondence between definably closed subsets of $B$ containing $A$ and closed subgroups $H$ of ${\mathcal{G}}=Aut(B/A)$. Namely $C$ corresponds to $H_{C}=\{\sigma\in{\mathcal{G}}:\sigma$ fixes pointwise C}C\}, $H$ corresponds to $C_{H}$ the set of elements of $B$ fixed by $H$, and $C_{H_{C}}=C$ and $H_{C_{H}}=H$.
Finally in this section we make the connection with works of Poizat referred to above. In the context of $T$ totally transcendental, Poizat defines $A_{min}$, the minimal closure of $A$ to be the intersection of all $f(M_{A})$ as $f$ ranges over $A$-elementary embeddings of $M_{A}$ into itself. The notation $A_{min}$ comes from [11]. In [13] Poizat uses the notation $Cl(A)$ for $A_{min}$ where an equivalent definition of $A_{min}$ is given, in terms of atomic algebraicity. In this paper we will use the notation $mcl(A)$.
Lemma 2.14.
Let $mcl(A)$ be the minimal closure of $A$ in $M_{A}$.
  1. 1.
    $mcl(A)$ is the greatest (under inclusion) $B$ with $A\subseteq B\subseteq M_{A}$ such that $B$ is normal and minimal (over $A$ in $M_{A}$).
  2. 2.
    Any subset of $mcl(A)$ is minimal.
Proof.
  1. 1.
    Corollaire 7 of [13] (actually there are two Corollaire 7’s) say that $f(mcl(A))=mcl(A)$ for every $A$-automomorphism of $M_{A}$, so we have normality. Corollaire 5 of [13] says in particular that $mcl(A)$ is atomic over any subset of itself which contains $A$, so if $B\supseteq A$ is a proper subset of $mcl(A)$, we could not have that $B\leq_{TV}mcl(A)$.
    Now suppose that $B\supseteq A$ is normal and minimal (over $A$ in $M_{A}$). Let $f$ be an $A$-elementary embedding of $M_{A}$ into itself.
    As $B$ is normal, $f(B)\subseteq B$ and we clearly have $f(B)\leq_{TV}B$. By minimality of $B$, $f(B)=B$. Hence $B\subseteq f(M)$, so $B\subseteq mcl(A)$.
  2. 2.
    Let $B$ be a subset of $mcl(A)$. Suppose $B$ is not minimal. Then there is a $C\leq_{TV}B$ with $C$ properly contained in $B$. Let $b\in B\backslash C$, then as $M_{A}$ is still prime over any subset of $mcl(A)$ [11], $M_{A}$ is prime over $C$. So let $\phi(\bar{x})$ isolate $tp(b/C)$. By $C\leq_{TV}B$, there is a $c\in C$ realizing $\phi(\bar{x})$, so the formula $x=c$ is in this type. This contradicts $b\in B\backslash C$.
So bearing in mind the last lemma, Poizat’s results, especially from [13] give a lot of information about minimal normal sets.
3 The Picard-Vessiot closure
The ambient theory here is $DCF_{0}$ the theory of differentially closed fields of characteristic 0 in the language $L_{\partial}$ of rings with a derivation (so the ring language) together with symbol $\partial$. We may use just $L$ for the ring language. $DCF_{0}$ is complete, $\omega$-stable, with quantifier elimination in $L_{\partial}$. $DCF_{0}$ also has elimination of imaginaries so $DCF_{0}=DCF_{0}^{eq}$.
We work in a big saturated model $\mathcal{U}$} of $DCF_{0}$. $K$ will denote a “small” differential subfield. We will write $K^{diff}$ (for differential closure of $K$) for the prime model $M_{K}$ over $K$. We assume that $C_{K}$, the field of constants of $K$, is algebraically closed. This assumption has a number of consequences, for example $C_{K^{diff}}=C_{K}$.
We repeat material from Section 3 of [9] and Section 1.2 of [6]. See also [4], [5]. However as mentioned in the introduction we will follow the conventions from [14]. A (homogeneous) linear differential equation (in vector form) over $K$ is an equation of the form $\partial Y=AY$ where $Y$ is a $n\times 1$ column vector of indeterminates and $A$ is an $n\times n$ matrix over $K$. The solution set ${\mathcal{Y}}(K^{diff})$ of the equation in $K^{diff}$ (a collection of $n\times 1$ column vectors from $K^{diff}$) is an $n$-dimensional vector space over $C_{K}=C_{K^{diff}}$. The (differential) field $L$ generated over $K$ by the set of all these solutions, equivalently by a $C_{K}$-basis for the solution set, is a called the $PV$ (Picard-Vessiot) extension of $K$ for the given equation. Given a family $\mathcal{F}$} of linear $DE$’s over $K$ (say $\{\partial Y_{i}=A_{i}Y_{i}:i\in I\}$) by the $PV$-extension of $K$ for the family $\mathcal{F}$} we mean the differential field extension of $K$ generated by the solutions in $K^{diff}$ of all these equations. By definition a $PV$ extension $L$ of $K$ is a $PV$ extension for some such family $\mathcal{F}$}. We say that $L$ is of finite type if $\mathcal{F}$} is a single equation (equivalently finitely many equations). Note that a $PV$ extension $L$ of $K$ is normal and minimal over $K$ (in $K^{diff}$). Normality is immediate. For minimality: suppose $K\leq L_{1}\leq_{TV}L$. Then $L$ will contain a $C_{K}$-basis of the solution set (in $K^{diff}$) of each equation in $\mathcal{F}$}. So $L_{1}=L$.
Let us note in passing that the first order vector form formalism for linear differential equations subsumes and is equivalent to the formalism of higher order homogeneous linear differential equations of one variable. An order $n$ such equation over $K$ is of the form $y^{(n)}+a_{n-1}y^{(n-1)}+...+a_{1}y^{\prime}+a_{0}y$ where the $a_{i}\in K$. $K^{PV}$ is generated by solutions in $K^{diff}$ of such equations. Moreover given such an equation, one can form an equation in vector form over $K$, $\partial Y=AY$ where $Y$ is an $n\times 1$ column vector of indeterminates such that the solutions of the latter are precisely of the form $(y,\partial(y),...,\partial^{(n-1)}(y))^{t}$ for $y$ a solution of the order $n$ equation.
When $L$ is a $PV$ extension of $K$ for a single equation $\partial Y=AY$ then $Aut(L/K)$ (as differential fields) has the structure of an algebraic subgroup of $GL_{n}(C_{K})$ (where $n$ is the length of the column vector $Y$). In general $Aut(L/K)$ is an inverse limit of such groups, so a proalgebraic group with respect to the algebraically closed field $C_{K}$.
By $K^{PV}$ we mean the $PV$-extension of $K$ for the family of all linear $DE$’s over $K$. $K^{PV}$ is clearly the greatest $PV$ extension of $K$. (Magid writes $K_{1}$ for $K^{PV}$.)
As $K^{PV}$ is normal over $K$ in $K^{diff}$ we have, by Remark 2.10 that $K^{diff}$ is also a differential closure of $K^{PV}$. So we can continue the construction to obtain $K^{PV_{2}}=(K^{PV})^{PV}$ the diffferential field generated by all solutions in $K^{diff}$ of all linear differential equations over $K^{PV}$. Then $K^{PV_{2}}$ is also normal (and minimal) over $K$ in $K^{diff}$, whereas $K^{diff}$ is also a differential closure of $K^{PV_{2}}$. We continue this was to build $K^{PV_{n}}$ inside $K^{diff}$ for all $n$. We let $K^{PV_{\infty}}=\cup_{n}K^{PV_{n}}$. To summarise:
Lemma 3.1.
(i) Each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ is normal and minimal over $K$ inside $K^{diff}$, i.e. $K\subseteq K^{PV}\subseteq(K^{PV})^{PV}\subseteq...\subseteq K^{PV_{\infty}}\subseteq mcl(K)\subseteq K^{diff}$.
(ii) $K^{PV_{\infty}}$ is $PV$-closed (has no proper $PV$ extension) and is in the fact $PV$-closure of $K$, smallest $PV$-closed differentrial subfield of $K^{diff}$ containing $K$.
Magid [5] denotes $K^{PV_{\infty}}$ by $K_{\infty}$ and calls it the complete $PV$-closure of $K$ (although of course “$PV$-closure” is correct).
In any case Proposition 2.13 applies to each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ to give a full Galois correspondence between closed subgroups of the automorphism group over $K$ and intermediate differential subfields.
Finally let us discuss the structure of differential subfields $L$ in between $K$ and $K^{PV_{\infty}}$ which are normal over $K$ in $K^{diff}$. The main point is that any such $L$ is an “iterated” Picard-Vessiot extension of $K$.
Proposition 3.2.
Let $L$ be a differential field such that $K\leq L\leq K^{PV_{\infty}}$ which is normal (over $K$ in $K^{diff}$). Let $L_{n}=L\cap K^{PV_{n}}$ (and take $L_{0}=K$). Then
(i) Let $N_{L}\ =\{n<\omega:L_{n+1}\setminus L_{n}\neq\emptyset\}$. Then $N_{L}$ is an initial segment of $\omega$ (so all of $\omega$ or a finite proper initial segment).
(ii) Each $L_{n+1}$ is a Picard-Vessiot extension of $L_{n}$ (maybe trivial if $L_{n+1}=L_{n}$) and is normal and minimal over $K$ in $K^{diff}$
(iii) $L$ is also minimal (over $K$ in $K^{diff}$).
Proof.
For simplicity we will consider just the case of $L_{1}$ and $L_{2}$. The reader can easily generalize the proof. Note that $L_{1}$ is a $PV$ extension (maybe trivial) of $L_{0}=K$. Suppose that $L_{2}\setminus L_{1}$ is nonempty. So, by our earlier discussion, there is $b\in L_{2}\setminus L_{1}$ which is a solution of an equation $y^{(n)}+a_{n-1}y^{(n-1)}+...+a_{1}y^{\prime}+a_{0}y$ with the $a_{i}\in L_{1}$ and $n>1$. We immediately replace $b$ by ${\bar{b}}=(b,\partial(b),..,\partial^{(n-1)}(b))^{t}$ (which note is in $L_{2}$) which is a solution of an appropriate $\partial Y=AY$ over $K_{1}$. Consider $tp({\bar{b}}/K^{PV})$ which is isolated by a formula $\phi({\bar{x}},c)$ where $c$ is a tuple from $K^{PV}$ which we may assume to be a “canonical parameter” of the formula $\phi({\bar{x}},c)$. Now as $K^{PV}$ is normal in $K^{diff}$, $c$ is fixed by any automorphism of $K^{diff}$ which fixes $\bar{b}$}. By homogeneity $c\in dcl({\bar{b}})$ in particular $c\in L$ so $c\in L_{1}$. Note that by normality of $L$ all realizations of $\phi({\bar{x}},c)$ are in $L_{2}$, and are moreover contained in $V$, the set of solutions of $\partial Y=AY$ in $K^{diff}$ (contained in $L_{2}$). $V$ is an $n$-dimensional vector space over $C_{K}$, and the set $X$ of realizations of $\phi({\bar{x}},c)$ is an infinite set of realizations of a complete type over $K^{PV}$). Then, using the Zilber indecomposability theorem for example, the $C_{K}$-vector subspace $W$ of $V$ generated by $X\cup\{0\}$ is definable, over $C_{K},c$, so over $L_{1}$. Now $W$ is contained in $L_{2}$, and contains the original chosen element $b$. Also on very general grounds the differential field generated over $L_{1}$ by $W$ is a (finite type) $PV$ extension of $L_{1}$. We have shown that $L_{2}$ is a $PV$ extension of $L_{1}$. ∎
So by the Proposition above we can use Proposition 2.13 to strengthen Theorem 5 of [5] by both replacing “normal locally iterated $PV$ extension” by just “normal” and including a definition of “closed subgroup”.
Corollary 3.3.
Let $L$ be a differential subfield of $K^{PV_{\infty}}$ containing $K$ and suppose $L$ is normal (over $K$ in $K^{diff}$). Then there is a full Galois correspondence between intermediate differential fields (between $K$ and $L$) and closed subgroups of $Aut(L/K)$.
4 $Aut(K_{2}/K)$
Our aim to prove Proposition 1.4 in the introduction. So $K$ is again a differential field with algebraically closed field $C_{K}$ of constants. Using the notation now of Section 3, we take $L$ to be a normal differential subfield of $K^{PV_{\infty}}$ such that $L=L_{2}$ and $L_{1}\neq L_{2}$. Equivalently, $L$ is a normal differential subfield of $K^{PV_{2}}$ which is not contained in $K^{PV}$, equivalently, $L$ is not a $PV$-extension of $K$. By Proposition 3.2, $L_{1}$ is a (nontrivial) $PV$ extension of $K$ and $L_{2}$ is a $PV$ extension of $L_{1}$. (Moreover there is no differential subfield $F$ of $L_{2}$ properly containing $L_{1}$ which is a $PV$ extension of $K$.)
A special case is, of course, taking $L_{1}=K^{PV}$ and $L_{2}=K^{PV_{2}}$ which is the case appearing in the abstract.
We will now change notation to be consistent with the introduction. So we will write $K_{1}$ for $L_{1}$ and $K_{2}$ for $L_{2}$ and reserve the symbol $L$ for other things.
We let $G=Aut(K_{2}/K)$, $N=Aut(K_{2}/K_{1})$ and $H=Aut(K_{1}/K)$. So we have the exact sequence
$1\to N\xrightarrow{i}G\xrightarrow{\pi}H\to 1$
of abstract groups for now, where $i$ is the natural embedding.
[…]
Let us also remark that the group $G=Aut(K_{2}/K)$ above could not in any reasonable sense have the structure of an inverse limit of definable groups. In that case it would be a proalgebraic group over $C_{K}$ which is a extension of an pro-affine algebraic group by a pro-affine algebraic group, so would itself be pro-affine. But then $K_{2}$ would itself be a $PV$ extension of $K$, which it is not. In any case we would like to recover some additional (pro-)definable or proalgebraic structure on $G$. This is Proposition 1.4.
Let us fix $D$ a left/right coset of $N$ in $G$, So $D=\pi^{-1}({\sigma_{0}})$ for some $\sigma_{0}\in H$. Namely $D=\{\tau\in G=Aut(K_{2}/K):\tau|K_{1}=\sigma_{0}\}$. We will take $\sigma_{0}\neq id$. Let us fix a $PV$ extension $L$ of $K_{1}$ contained in $K_{2}$. Then for $\tau\in D$, $\tau(L)$ depends only on $\sigma_{0}$ because $L$ is generated by a solution to a linear ODE defined over $K_{1}$. So let $L^{\prime}=\tau(L)$ for some/any $\tau\in D$. Let $D_{L}=\{\tau|L:\tau\in D\}$. Note that as $D$ is a left coset $\tau N$ of $N$, we have a right action of $N$ on $D$ (multiplication in $G$), and likewise we have a left action of $N$ on $D$.
Remark 4.1.
(i) Restriction to $L$ gives a right (strictly transitive) action of $Aut(L/K_{1})$ on $D_{L}$.
(ii) Restriction to $L^{\prime}$ gives a left (strictly transitive) action of $Aut(L^{\prime}/K_{1})$ on $D_{L}$.
(iii) So we have a bi-torsor $(Aut(L^{\prime}/K_{1}),D_{L},Aut(L/K_{1}))$.
Proof.
Immediate. The bi-torsorial nature (left and right actions commute) is because everything is induced by multiplication in $G$ (which is associative).
Proposition 4.2.
The bitorsor in (iii) above is $K_{1}$-definable. More specifically there is a bijection $h:D_{L}\to C_{L}$ for some $K_{1}$-definable $C_{L}$, and $K_{1}$-definable left action of $H_{L^{\prime}}$ on $C_{L}$, and $K_{1}$-definable right action of $H_{L}$ on $C_{L}$, such that $(h_{L^{\prime}},h_{L},h)$ give an isomorphism of bitorsors between
$(Aut(L^{\prime}/K_{1}),D_{L},Aut(L/K_{1}))$ and $(H_{L^{\prime}},D_{L},H_{L})$.
We will make use of the analysis from Sections 3 and 4 of [1] which we now summarize to fix notation (before beginning the proof of Proposition 4.2).
Let us suppose that our definitions of $G_{L}$ and $H_{L}$ earlier in this section depended upon choosing a linear DEDE $\partial Y=AY$ and a fundamental matrix of solutions $b$ such that $L=K(b)$. So $b$ is a single solution of the logarithmic differential equation $\partial Z=AZ$ where $Z$ is a matrix of indeterminates ranging over $GL_{n}$. Now let ${\mathcal{Y}}_{A}$ denote the solution set of $\partial Z=AZ$ in $K^{diff}$ (so ${\mathcal{Y}}_{A}\subseteq GL_{n}(L)$). As defined earlier, $h_{L}$ denotes the map from $Aut(L/K_{1})$ to $GL_{n}(K^{diff})$ taking $\sigma$ to $\sigma(b)b^{-1}$ (multiplication and inversion in the sense of $GL_{n}$). Note that $\sigma\in Aut(L/K_{1})$ is determined by its action on ${\mathcal{Y}}_{A}$, As in Lemma 3.3 of [1] $\rho$ does not depend on the choice of $b\in{\mathcal{Y}}_{A}$, and moreover establishes the isomorphism $h_{L}$ between $Aut(L/K_{1})$ acting on ${\mathcal{Y}}_{A}$ and a differential algebraic subgroup $H_{A}^{+}$ of $GL_{n}(K^{diff})$ acting on ${\mathcal{Y}}_{A}$ by left multiplication in $GL_{n}(K^{diff})$. $H_{A}^{+}$ is what we called $H_{L}$ earlier but we use the new notation as it is consistent with [1] and also emphazises the dependence on $A$. $H_{A}^{+}$ is the “intrinsic” differential Galois group of $L$ over $K_{1}$.
Now ${\mathcal{Y}}_{A}$ is a union of $H_{A}^{+}$ orbits, equivalently (in $GL_{n}(K^{diff})$) a union of right cosets of $H_{A}^{+}$. For some $m$ there is an $A$-definable subset $O_{A}$ of $(C_{K})^{m}$ and an $A$-definable surjective function $f_{A}:{\mathcal{Y}}_{A}\to O_{A}$ inducing a bijection between ${\mathcal{Y}}_{A}/H_{A}^{+}$ and $O_{A}$. By construction the complete types over $K_{1}$ of fundamental matrices of solutions for $\partial Y=AY$ correspond to the fibres $f_{A}^{-1}(a^{\prime})$ for $a^{\prime}\in O_{A}$. Let $a=f_{A}(b)$ where $b$ was our above choice of fundamental matrix of solutions. The algebraic subgroup $G_{L}$ of $GL_{n}(C_{K})$ constructed earlier depends only on $A$ and $a$ and is precisely $\{b_{2}^{-1}\cdot b_{1}:b_{1},b_{2}\in({\mathcal{Y}}_{A})_{a}$}. In [1] it is denoted $(H_{A})_{a}$.
Let $\sigma_{0}(A)=B$. So $L^{\prime}$ is the $PV$ extension of $K_{1}$ for the linear DEDE $\partial Y=BY$. We can define ${\mathcal{Y}}_{B}$, $(H_{B})^{+}=H_{L^{\prime}}$, $O_{B}$ and $f_{B}:{\mathcal{Y}}_{B}\to O_{B}$ as before.
Just for the record:
Remark 4.3.
For any $\tau\in D_{L}$ we have:
(i) $\tau({\mathcal{Y}}_{A})={\mathcal{Y}}_{B}$.
(ii) $\tau(f_{A})=f_{B}$.
(iii) $O_{A}=O_{B}$ and is fixed pointwise by $\tau$.
(iv) $f_{B}(\tau(b))=a=f_{A}(b)$.
(v) $\tau$ is determined by $\tau(b)$.
(vi) $({\mathcal{Y}}_{B})_{a}$ is precisely the set of $\tau^{\prime}(b)$ as $\tau^{\prime}$ ranges over $D_{L}$
Proof. (i) to (iv) are clear. For (v), every element $b_{1}$ of ${\mathcal{Y}}_{A}$ is of the form $b\cdot c$ for some $c\in GL_{n}(C_{K})$ so $\tau(b_{1})=\tau(b\cdot c)=\tau(b)\cdot c$.
(vi): Notice that $tp(b/K_{1})$ is isolated by $y\in({\mathcal{Y}}_{A})_{a}$”, hence $\sigma_{0}(tp(b/K_{1}))$ is isolated by $y\in({\mathcal{Y}}_{B})_{a}$” so $\tau(b)$ in $({\mathcal{Y}}_{B})_{a}$ and for any $d\in({\mathcal{Y}}_{B})_{a}$ there is $\tau^{\prime}\in D_{L}$ such that $\tau^{\prime}(b)=d$. ∎
Proof of Proposition 4.2. Recall our fixed $b\in{\mathcal{Y}}_{B}$. Let $C_{L}=\{\tau(b)\cdot b^{-1}:\tau\in D_{L}\}$. By Remark 4.3 (v) the map $h$ taking $\tau\in D_{L}$ to $\tau(b)\cdot b^{-1}$ is a bijection between $D_{L}$ and $C_{L}$.
Claim 1. $C_{L}$ and $h:D_{L}\to C_{L}$ do not depend on the choice of $b\in{\mathcal{Y}}_{A}$.
Proof of claim. Let $b_{1}\in{\mathcal{Y}}_{A}$. So $b_{1}=b\cdot c$ for some $c\in GL_{n}(C_{K})$. So for any $\tau\in D_{L}$,
$\tau(b_{1})\cdot b_{1}^{-1}=\tau(b\cdot c)\cdot(b\cdot c)^{-1}=\tau(b)\cdot c\cdot c^{-1}\cdot b^{-1}$. ∎
Recall again that the isomorphism $h_{L}:Aut(L/K_{1})\to H_{L}=(H_{A})^{+}$ taking $\sigma$ to $\sigma(b)\cdot b^{-1}$ does notdepend on the choice of $b\in{\mathcal{Y}}_{A}$.
Claim 2. Let $\sigma\in Aut(L/K_{1})$ and $\tau\in D_{L}$, then $h(\tau\sigma)=h(\tau)\cdot h_{L}(\sigma)$. (where $\tau\sigma$ is multiplication in $G$ and $\cdot$ is multiplication in $GL_{n}(K^{diff})$.
Proof of claim. So fix $\tau$ and $\sigma$. By Claim 1, $h(\tau)=\tau(\sigma(b))\cdot\sigma(b)^{-1}$ as $\sigma(b)\in{\mathcal{Y}}_{A}$.
Hence $h(\tau)\cdot h_{L}(\sigma)=\tau(\sigma(b))\cdot\sigma(b)^{-1}\cdot\sigma(b)\cdot b^{-1}=\tau\sigma(b)\cdot b^{-1}=h(\tau\sigma)$. ∎
Claim 2 shows the definability (over $K_{1}$) of the right action of $Aut(L/K_{1})$ on $D_{L}$. A similar proof gives definability of the right action of $Aut(L^{\prime}/K_{1})$ on $D_{L}$. This completes the proof of Proposition 4.2. ∎.
Corollary 4.4.
Fix $\tau\in G$. Then conjugation by $\tau$ induces a definable isomorphism between the definable groups $H_{A}^{+}$ and $H_{B}^{+}$.
Proof.
For $\sigma\in Aut(L/K_{1})$, $\tau\sigma\tau^{-1}$ is the unique $\sigma^{\prime}\in Aut(L^{\prime}/K_{1}$ such that $\tau\sigma=\sigma^{\prime}\tau$. Identifying via $h_{L}$, $D_{L}$ with $C_{L}$, and $Aut(L/K_{1})$, $Aut(L^{\prime}/K_{1})$ with $H_{A}^{+}$, $H_{B}^{+}$ respectively, Proposition 4.2 yields that maps $H_{A}^{+}$ to $C_{L}$ taking taking $\sigma$ to $\tau\sigma$, and from $H_{B}^{+}$ to $C_{L}$ taking $\sigma^{\prime}$ to $\sigma^{\prime}\tau$ are definable. This suffices. ∎
Corollary 4.5.
[…]
Proof.
[…]
Remark 4.6.
(i) In fact in Proposition 4.2, and Corollaries 4.3 and 4.4 we do not make any use of the assumption that $K_{1}$ is a $PV$-extension. We use nothing about the Picard-Vessiot extension $K_{1}$ of $K$. So the content is as follows: Suppose $C_{K}$ is algebraically closed. Let $K\leq K_{1}\leq K_{2}\leq K^{diff}$ be such that $K_{1}$ and $K_{2}$ are normal over $K$ in $K^{diff}$ and $K_{2}$ is a Picard-Vessiot extension of $K_{1}$. The conjugation by any given element of $G=Aut(K_{2}/K)$ is a proalgebraic automorphism of $N=Aut(K_{2}/K_{2})$ where the latter is considered as the group of $C_{K}$-points of a proalgebraic group over $C_{K}$.
(ii) However the next Corollary does make use of the Picard-Vessiot extension $K_{1}$ of $K$.
Corollary 4.7.
Suppose $C_{K}$ is algebraically closed, $K_{1}$ is a $PV$ extension of $K$ and $K_{2}$ a $PV$ extension of $K_{1}$ and $K_{2}$ is normal over $K$ in $K^{diff}$. Suppose moreover that $G=Aut(K_{2}/K)$ is a semidirect product $N\rtimes H$ of $N=Aut(K_{2}/K_{2})$ and $H=Aut(K_{1}/K)$. Then for any $g\in N$, left translation by $g$ is a proalgebraic morphism in $N\times H$ viewed as a proalgebraic variety.
Proof.
So as both $N$ and $H$ have the structure of proalgebraic varieties (over $C_{K}$), so does the Cartesian product $N\times H$. Fix $g\in G$ which we write as $(n_{1},h_{1})\in N\times H$. Then left multiplication by $g$ in the group $G$ takes $(n_{2},h_{2})$ to $(n_{1}n_{2}^{h_{1}},h_{1}h_{2})$. As by Corollary 4.5, conjugation by the fixed element $h_{1}\in G$ is a proalgebraic automorphism of $N$ and both multiplication in $N$ and in $H$ are proalgebraic, it follows that left multiplication by $g$ is a proalgebraic morphism from $N\times H$ to itself.